one of the two complementary angles measures 30 degrees more than three times the other. Find the measure of each angle
Answer by stanbon(75887)

one of the two complementary angles measures 30 degrees more than three times the other. Find the measure of each angle
------
1st angle: x
2nd angle: 90-x
-------------------
Equation:
x = 3(90-x)+30
-----
x = 270 - 3x + 30
4x = 300
x = 75 degrees
90-x = 15 degrees
=========================
Cheers,
Stan H.